OpenStudy (anonymous):

48÷2(9+3)= ? Is the answer going to be 2 or 288? And why?

sam (.sam.):

PEMDAS 288

OpenStudy (anonymous):

But Isn't Multiplication in the same priority of division, so why?

OpenStudy (anonymous):

48 / 18 + 6 6+6 = 12

OpenStudy (anonymous):

The priority is B-O-D-M-A-S.

OpenStudy (anonymous):

BODMAS? I'm unfamiliar with that one, woudl you care to explain?

sam (.sam.):

In the acronym PEMDAS, the M and D operations are at the same priority and the A and S are at the same priority. Arithmetic operations at the same priority are evaluated left to right. So for example, 2 + 5 - 3 is the same as (2 + 5) - 3 = 7 - 3 = 4, while 2 - 5 + 3 is the same as (2 - 5) + 3 = -3 + 3 = 0.

OpenStudy (anonymous):

B Brackets first O Orders (ie Powers and Square Roots, etc.) DM Division and Multiplication (left-to-right) AS Addition and Subtraction (left-to-right) source http://www.mathsisfun.com/operation-order-bodmas.html

OpenStudy (anonymous):

Oh alright, thanks guy :D